# BEAT: Balanced Frequency Adaptive Tuning for Long-Term Time-Series Forecasting

### v1

@@ Abstract
Machine Learning, ICML
[…]
1 Introduction
Figure 1: Comparison of the learning process when adjusted using the proposed BEAT approach. Training is conducted on the ETTh2 dataset with three levels of wavelet decomposition. The symbols $\blacktriangle$ and $\bullet$ mark the iteration where each training curve converged. BEAT dynamically reallocates priorities to ensure balanced training and improved performance against the original one.
Long-term time-series forecasting plays a pivotal role in various real-world applications, such as weather prediction (Zhang et al., 2022), electric load forecasting (Zhou et al., 2021; Gasparin et al., 2022), traffic flow analysis (Jin et al., 2021), and financial market modeling (Lai et al., 2018; Tang et al., 2022). These tasks demand accurate prediction over extended temporal horizons. In recent years, deep learning methods for time-series analysis have achieved significant advancements, characterized by the development of various architectures, including CNN-based (Liu et al., 2022; Wang et al., 2023; Luo & Wang, 2024), MLP-based (Oreshkin et al., 2019; Zeng et al., 2023; Wang et al., 2024a), Transformer-based (Wu et al., 2021; Liu et al., 2021, 2024), and LLM-based (Jin et al., 2024; Jia et al., 2024) approaches. Compared with temporal-domain approaches, frequency-domain methods (Wu et al., 2021, 2023; Zhou et al., 2022b) transform time-series data into the frequency domain, enabling the effective capture of large-scale periodic patterns, reducing sequence dependency lengths, and enhancing robustness through inherent denoising effects.
Despite these advantages, frequency-domain analysis typically decomposes time-series data into multiple sub-series, such as high-frequency, low-frequency, and potentially additional band-pass segments, each representing a distinct temporal scale. High-frequency sub-series often capture short-term fluctuations or noise, while low-frequency sub-series reflect long-term trends or stable periodic patterns. Existing frequency-domain approaches, such as WPMixer (Murad et al., 2025), simply train all frequency sub-series simultaneously under a unified loss function. However, as illustrated in Figure 1, the following challenges are observed: (1) the high-frequency component of the model tends to converge more rapidly, while the low-frequency component of the model requires more time to be adequately learned; (2) early stopping may result in underfitting of the low-frequency components, even if the high-frequency components are well-learned, whereas prolonged training risks overfitting the high-frequency components. This asynchronous learning pace across different frequencies renders a unified loss function insufficient to optimize all sub-series simultaneously, ultimately leading to suboptimal overall performance. Therefore a critical challenge lies in dynamically evaluating the learning progress of each frequency-specific sub-series and introducing timely adjustments to ensure balanced training for both high-frequency and low-frequency components.
@@ 1 Introduction
2 Related Work
2.1 Temporal-domain Time-Series Forecasting
@@
CNN-based methods. Convolutional neural networks (He et al., 2016; Gu et al., 2018) have gained prominence for their effectiveness in capturing local features and recognizing patterns for many tasks (Jiang et al., 2024a; Li et al., 2022). SCINet (Liu et al., 2022) employs standard convolutions with a hierarchical downsample-convolve-interact framework to capture dynamic temporal dependencies across varying temporal resolutions of time series data. MICN (Wang et al., 2023) integrates different convolutional kernels to model temporal correlations from both local and global perspectives. ModernTCN (Luo & Wang, 2024) enhances traditional TCN (Bai et al., 2018)by employing DWConv and ConvFFN to capture cross-time and cross-variable dependencies independently.
@@
MLP-based methods. Drawing inspiration that outputs depend linearly on their historical values, Multi-Layer Perceptrons (MLP) have emerged as a favored structure for time series modeling. These approaches handle multivariate data as independent univariate sequences and have shown promising results (Zeng et al., 2023). N-BEATS (Oreshkin et al., 2019), as a purely MLP-driven deep learning model for time series, utilizes deep stacks of fully-connected layers with dual residual branches - one for backcasting and the other for forecasting - per layer. TimeMixer (Wang et al., 2024a) proposing an MLP-based multiscale mixing framework posits that time series exhibit varying patterns across different sampling scales.
@@
Transformer-based methods. Motivated by the remarkable success of the Transformer architecture  (Vaswani, 2017) in sequence-to-sequence tasks (Li et al., 2023; Jiang et al., 2024b), prevailing it to the domain of time series modeling  (Zhou et al., 2022b, 2021; Wu et al., 2021). Autoformer (Wu et al., 2021) pioneers an Auto-Correlation Mechanism. A time delay module is introduced to aggregate similar sub-series from inherent periods instead of focusing on scattered points. Pyraformer (Liu et al., 2021) constructs a multi-resolution C-ary tree and proposes a Pyramidal Attention Mechanism, enabling it to capture both short- and long-temporal dependencies in linear time and space. iTransformer (Liu et al., 2024) further expands the receptive field by tokenizing the entire time series to capture inter-series dependencies.
2.2 Frequency-domain Time-Series Forecasting
@@
Due to the presence of both high- and low-frequency components, purely time-domain methods may struggle to manage all frequency bands effectively in long-range scenarios compared with frequency-domain approaches. The ability to decompose signals into sub-series across different frequency ranges makes them more suitable for long-horizon forecasting.
@@
As a common practice, TimesNet (Wu et al., 2023) applies FFT (Brigham, 1988) to isolate dominant frequencies—those with the highest amplitudes—and rearranges 1D time series into 2D representations based on identified periods. Autoformer (Wu et al., 2021) introduces an Auto-Correlation mechanism alongside efficient FFT to treat data as a real discrete-time signal and capture series-wise correlations. FiLM (Zhou et al., 2022a) employs Frequency Enhanced Layers (FEL), merging Fourier analysis with low-rank approximation to retain key low-frequency components and top eigenspaces, thereby reducing noise and expediting training. FITS (Xu et al., 2023) uses a low-pass filter (LPF) to remove high-frequency elements beyond a designated cutoff, shrinking model size while maintaining essential information. From the opposite idea, FEDformer (Zhou et al., 2022b) argues that focusing solely on low-frequency components can miss critical variations in time series data. Consequently, to capture a global perspective, the FEDformer randomly selects a fixed number of Fourier components—encompassing both high- and low-frequency ranges. EV-FGN (Yi et al., 2022) uses a two-dimensional DFT on the spatial-temporal plane of embeddings, combined with graph convolutions to jointly model spatial-temporal dependencies in the frequency domain. FreTS (Yi et al., 2024) leverages the DFT to transform data into frequency-domain spectra and introduces frequency-domain MLPs designed for complex inputs by separately modeling real and imaginary parts. FCVAE (Wang et al., 2024c) incorporates both global and local frequency components into the condition of a Conditional Variational Autoencoder (CVAE). TSLANet (Eldele et al., 2024) proposes a lightweight Adaptive Spectral Block (ASB) that replaces self-attention using Fourier-based global and local filters. The recent WP-Mixer (Murad et al., 2025)employs multi-level wavelet decomposition, combining patching and patch mixer for local and global information respectively, successfully handling complex data behaviors such as sudden spikes and drops. However, WPMixer faces key issues: it simultaneously trains all frequency sub-series under a unified loss function, with high-frequency components learning quickly while low-frequency components lag, risking underfitting and overfitting due to asynchronous learning. In contrast, our model BEAT handles this by decomposing the target into frequency-specific sub-series and evaluating their convergence statuses. It adjusts gradients in real time to achieve synchronization of the learning progress across frequencies, overcomes uniform loss limitations, and improves model efficiency.
@@
Figure 2: Overall architecture of the proposed BEAT. (a) and (b) are the illustrations of the wavelet decomposition and reconstruction, utilizing the multi-level discrete wavelet transform and the inverse operation. (c) demonstrates the process of monitoring the discrepancy ratio for each frequency. (d) presents the process of calculating the gradient modulation coefficients which are applied to adjust the gradients being back-propagated. Best viewed in color and zoom in.
3 Methodology
3.1 Task Formulation
@@
3.2 Overall Architecture
As shown in Figure 2, the overall architecture of the proposed BEAT is presented. For the input historical time series sequence $X$, the Reversible Instance Normalization (RevIN) (Kim et al., 2021) is employed for handling the varied mean and variation in the time-series data. Then the normalized data $X_{norm}$ is transposed and decomposed by the multi-level wavelet decomposition (Mallat, 1989; Daubechies, 1992; Wang et al., 2018) as approximation $X_{A}$ and detail coefficients $X_{D_{i}},i\in\{1,2,\dots,f\}$ series corresponding to $m$ frequency levels from low to high.
@@ 3.2 Overall Architecture
To demonstrate our proposed BEAT approach, specifically tailored for frequency-domain methods, we utilize WPMixer (Murad et al., 2025) as an instantiation. The approximation coefficient series at level $f$ is preserved, ensuring a concise and efficient representation.
@@
3.3 Frequency-Specific Monitor
The Frequency-Specific Monitor (FSM) is designed to track the training status of the network for each frequency, supporting the dynamical adjustment of the learning pace for these networks inspired by (Peng et al., 2022). To monitor the learning status for each frequency, it is necessary to compare their prediction with corresponding targets. To this end, we take advantage of the ground-truth sequence $Y$ and utilize the wavelet decomposition to transform $Y$ in the temporal domain as the approximation coefficient series $Y_{A}$ and detail coefficient series $Y_{D_{i}}$ where $i\in\{1,2,\dots,f\}$ and $f$ is the wavelet decomposition level. The wavelet decomposition process can be formulated as:
@@ 3.3 Frequency-Specific Monitor
3.4 Dynamical Gradient Balancer
@@ 3.5 Training and Inference
Table 1: Dataset descriptions of datasets utilized for evaluation.
Dataset # VAR Size FRQ Domain
Weather 21 (36792, 5271, 10540) 10 min Weather
ETTh1, ETTh2 7 (8545, 2881, 2881) 1 hour Electricity
ETTm1, ETTm2 7 (34465, 11521, 11521) 15 min Electricity
Traffic 862 (12185, 1757, 3509) 1 hour Transportation
ECL 321 (18317, 2633, 5261) 1 hour Electricity
@@
4 Experiments
4.1 Experimental Settings
Table 2: Comparison with state-of-the-art methods on the test sets of seven challenging datasets for long-term time-series forecasting. The historical time step is set as 96 for fairness. The prediction length varies in the range of $\{96,192,336,720\}$. Counts of ranking first for each metric across all methods are reported. For each metric, the best performance is in bold while the second is underlined.
Ours 2025 2024 2024 2023 2023 2023 2023 2023
@@ 4.1 Experimental Settings
[…]
1st Count: 28 30 0 0 1 0 7 6 0 0 0 0 0 0 0 0 0 0
Datasets.
We benchmark our method over seven challenging datasets in long-term series forecasting. (1) Weather dataset [^1] https://www.bgc-jena.mpg.de/wetter/ is collected every 10 minutes throughout 2020, containing 21 meteorological indicators such as air temperature and humidity. (2) ETT dataset (Zhou et al., 2021) contains data collected from electricity transformers, including load and oil temperature measurements recorded every 15 minutes from July 2016 to July 2018. There are four sub-datasets in ETT, including ETTh1, ETTh2, ETTm1, and ETTm2. (3) Traffic dataset [^2] https://pems.dot.ca.gov/ comprises hourly data from the California Department of Transportation, detailing road occupancy rates measured by various sensors on San Francisco Bay Area freeways. (4) ECL dataset (Li et al., 2019) records hourly for the electricity consumption data from 321 clients. Table 1 provides data statistics of these datasets. Size denotes the number of time points in the (Train, Validation, Test) sets, respectively.
Baselines.
There are nine state-of-the-art baseline methods for long-term time series forecasting are compared, including WPMixer (Murad et al., 2025), TimeMixer (Wang et al., 2024a), iTransformer (Liu et al., 2024), TSMixer (Ekambaram et al., 2023), PatchTST (Nie et al., 2023), Crossformer (Zhang & Yan, 2023), TiDE (Das et al., 2023), and DLinear (Zeng et al., 2023).
Implementation Details.
We develop BEAT using Pytorch (Paszke et al., 2019) based on the TSLib (Wang et al., 2024b) codebase. Following the previous method (Ekambaram et al., 2023), we set the historical time step $T$ as 96 and follow the unified setting for a fair comparison for all experiments. The prediction length $K$ varies within the range of $\{96,192,336,720\}$. For each dataset, all experiments are conducted on the same training set and evaluated on the same test set. A single NVIDIA RTX 4090 with 24GB memory is used for all experiments.
@@
4.2 Long-Term Time Series Forecasting Results
@@
Table 3: Ablation study of designed components in BEAT. FSM and DGB represent the proposed Frequency-Specific Monitor and Dynamical Gradient Balancer. The average performance over four prediction lengths including $\{96,192,336,720\}$ are reported for each dataset’s test set. Best in bold, second underlined.
Index FSM DGB Weather ETTh1
MSE MAE MSE MAE
(1) 0.243 0.269 0.422 0.423
(2) 0.239 0.263 0.415 0.419
4.3 Ablation Study
We conduct extensive ablation studies on the test sets of the Weather and ETTh1 datasets to evaluate the effectiveness of the proposed BEAT framework. To be specific, we examine BEAT from the aspect of technical designs, wavelet type, discrepancy calculation design, and gradient modulation design in the following.
Table 4: Ablation study of the design choice of the wavelet type. The average performance over four prediction lengths including $\{96,192,336,720\}$ are reported for each dataset’s test set. Best in bold, second underlined.
@@ 4.3 Ablation Study
Table 5: Ablation study of levels of decomposition in the wavelet transform. The average performance over four prediction lengths including $\{96,192,336,720\}$ are reported for each dataset’s test set. Best in bold, second underlined.
Index Levels of Decomposition $f$ Weather ETTh1
@@
Table 6: Ablation study of the design for discrepancy calculation. The average performance over four prediction lengths including $\{96,192,336,720\}$ are reported for each dataset’s test set. Best in bold, second underlined.
Index Discrepancy Calculation Weather ETTh1
MSE MAE MSE MAE
(1) RMSE 0.244 0.268 0.421 0.425
(2) R-Squared 0.241 0.265 0.418 0.423
(3) MSE 0.239 0.263 0.415 0.419
@@
Table 7: Ablation study of the design for learning process modulation. The average performance over four prediction lengths including $\{96,192,336,720\}$ are reported for each dataset’s test set. Best in bold, second underlined.
Index Modulation Weather ETTh1
MSE MAE MSE MAE
(1) Loss 0.242 0.268 0.420 0.421
(2) Gradient 0.239 0.263 0.415 0.419
@@
To evaluate the effect of the discrepancy calculation metric used in Equations 2 and 3, we conduct experiments on the metric for measuring the discrepancy between the predicted coefficients and the ground-truth ones, reflecting the learning process of networks for different frequencies. As shown in Table 6, using MSE for calculating the discrepancy can achieve the best performance against other metrics including RMSE, R-Squared, and MAE. MSE outperforms other metrics because by squaring the errors, it disproportionately penalizes larger discrepancies, leading to a more accurate measurement.

### v2

@@ Abstract
Introduction
Long-term time-series forecasting plays a pivotal role in various real-world applications, such as weather prediction (Zhang et al. 2022), electric load forecasting (Zhou et al. 2021; Gasparin, Lukovic, and Alippi 2022), traffic flow analysis (Jin et al. 2021), and financial market modeling (Lai et al. 2018; Tang et al. 2022). These tasks demand accurate prediction over extended temporal horizons. In recent years, deep learning methods for time-series analysis have achieved significant advancements, characterized by the development of various architectures, including CNN-based (Liu et al. 2022; Wang et al. 2023; Luo and Wang 2024), MLP-based (Oreshkin et al. 2019; Zeng et al. 2023; Wang et al. 2024a), Transformer-based (Wu et al. 2021; Liu et al. 2021, 2024), and LLM-based (Jin et al. 2024; Jia et al. 2024) approaches. Compared with temporal-domain approaches, frequency-domain methods (Wu et al. 2021, 2023; Zhou et al. 2022b) transform time-series data into the frequency domain, enabling the effective capture of large-scale periodic patterns, reducing sequence dependency lengths, and enhancing robustness through inherent denoising effects.
Despite these advantages, frequency-domain analysis typically decomposes time-series data into multiple sub-series, such as high-frequency, low-frequency, and potentially additional band-pass segments, each representing a distinct temporal scale. High-frequency sub-series often capture short-term fluctuations or noise, while low-frequency sub-series reflect long-term trends or stable periodic patterns. Existing frequency-domain approaches, such as WPMixer (Murad, Aktukmak, and Yilmaz 2025), simply train all frequency sub-series simultaneously under a unified loss function. However, the following challenges are observed: (1) the high-frequency component of the model tends to converge more rapidly, while the low-frequency component of the model requires more time to be adequately learned; (2) early stopping may result in underfitting of the low-frequency components, even if the high-frequency components are well-learned, whereas prolonged training risks overfitting the high-frequency components. This asynchronous learning pace across different frequencies renders a unified loss function insufficient to optimize all sub-series simultaneously, ultimately leading to suboptimal overall performance. Therefore a critical challenge lies in dynamically evaluating the learning progress of each frequency-specific sub-series and introducing timely adjustments to ensure balanced training for both high-frequency and low-frequency components.
@@ Introduction
Related Work
Temporal-domain Time-Series Forecasting
@@
CNN-based methods. Convolutional neural networks (He et al. 2016; Gu et al. 2018) have gained prominence for their effectiveness in capturing local features and recognizing patterns for many tasks (Jiang et al. 2024a; Li et al. 2022). SCINet (Liu et al. 2022) employs standard convolutions with a hierarchical downsample-convolve-interact framework to capture dynamic temporal dependencies across varying temporal resolutions of time series data. MICN (Wang et al. 2023) integrates different convolutional kernels to model temporal correlations from both local and global perspectives. ModernTCN (Luo and Wang 2024) enhances traditional TCN (Bai, Kolter, and Koltun 2018)by employing DWConv and ConvFFN to capture cross-time and cross-variable dependencies independently.
@@
MLP-based methods. Drawing inspiration that outputs depend linearly on their historical values, Multi-Layer Perceptrons (MLP) have emerged as a favored structure for time series modeling. These approaches handle multivariate data as independent univariate sequences and have shown promising results (Zeng et al. 2023). N-BEATS (Oreshkin et al. 2019), as a purely MLP-driven deep learning model for time series, utilizes deep stacks of linear layers with dual residual branches. TimeMixer (Wang et al. 2024a) proposing an MLP-based multiscale mixing framework posits that time series exhibit varying patterns across different sampling scales.
@@
Transformer-based methods. Motivated by the remarkable success of the Transformer architecture  (Vaswani 2017) in sequence-to-sequence tasks (Li et al. 2023; Jiang et al. 2024b), prevailing it to the domain of time series modeling  (Zhou et al. 2022b, 2021; Wu et al. 2021). Autoformer (Wu et al. 2021) pioneers an Auto-Correlation Mechanism. A time delay module is introduced to aggregate similar sub-series from inherent periods instead of focusing on scattered points. Pyraformer (Liu et al. 2021) constructs a multi-resolution C-ary tree and proposes a Pyramidal Attention Mechanism, enabling it to capture both short- and long-temporal dependencies in linear time and space. iTransformer (Liu et al. 2024) further expands the receptive field by tokenizing the entire time series to capture inter-series dependencies.
Frequency-domain Time-Series Forecasting
@@
Due to the presence of both high- and low-frequency components, purely time-domain methods may struggle to manage all frequency bands effectively in long-range scenarios compared with frequency-domain approaches. The ability to decompose signals into sub-series across different frequency ranges makes them suitable for long-horizon forecasting.
@@
As a common practice, TimesNet (Wu et al. 2023) applies FFT (Brigham 1988) to isolate dominant frequencies—those with the highest amplitudes—and rearranges 1D time series into 2D representations based on identified periods. Autoformer (Wu et al. 2021) introduces an Auto-Correlation mechanism alongside efficient FFT to treat data as a real discrete-time signal and capture series-wise correlations. FiLM (Zhou et al. 2022a) employs Frequency Enhanced Layers (FEL), merging Fourier analysis with low-rank approximation to retain key low-frequency components and top eigenspaces, thereby reducing noise and expediting training. FITS (Xu, Zeng, and Xu 2023) uses a low-pass filter (LPF) to remove high-frequency elements beyond a designated cutoff, shrinking model size while maintaining essential information. From the opposite idea, FEDformer (Zhou et al. 2022b) argues that focusing solely on low-frequency components can miss critical variations in time series data. Consequently, to capture a global perspective, the FEDformer randomly selects a fixed number of Fourier components, encompassing both high- and low-frequency ranges.
EV-FGN (Yi et al. 2022) uses a two-dimensional DFT on the spatial-temporal plane of embeddings, combined with graph convolutions to jointly model spatial-temporal dependencies in the frequency domain. FreTS (Yi et al. 2024) leverages the DFT to transform data into frequency-domain spectra and introduces frequency-domain MLPs designed for complex inputs by separately modeling real and imaginary parts. FCVAE (Wang et al. 2024c) incorporates both global and local frequency components into the condition of a Conditional Variational Autoencoder (CVAE). TSLANet (Eldele et al. 2024) proposes a lightweight Adaptive Spectral Block (ASB) that replaces self-attention using Fourier-based global and local filters.
The recent WP-Mixer (Murad, Aktukmak, and Yilmaz 2025) employs multi-level wavelet decomposition, combining patching and patch mixer for local and global information respectively, successfully handling complex data behaviors such as sudden spikes and drops. However, WPMixer faces key issues: it simultaneously trains all frequency sub-series under a unified loss function, with high-frequency components learning quickly while low-frequency components lag, risking underfitting and overfitting due to asynchronous learning. In contrast, our model BEAT handles this by decomposing the target into frequency-specific sub-series and evaluating their convergence statuses. It adjusts gradients in real time to achieve synchronization of the learning progress across frequencies, overcomes uniform loss limitations, and improves model efficiency.
@@
Figure 1: Overall architecture of the proposed BEAT. (a) and (b) are the illustrations of the wavelet decomposition and reconstruction, utilizing the multi-level discrete wavelet transform and the inverse operation. (c) demonstrates the process of monitoring the discrepancy ratio for each frequency. (d) presents the process of calculating the gradient modulation coefficients, which are applied to adjust the gradients being back-propagated. Best viewed in color and zoom in.
Methodology
Task Formulation
@@
Overall Architecture
As shown in Figure 1, the overall architecture of the proposed BEAT is presented. For the input historical time series sequence $X$, the Reversible Instance Normalization (RevIN) (Kim et al. 2021) is employed for handling the varied mean and variation in the time-series data. Then the normalized data $X_{norm}$ is transposed and decomposed by the multi-level wavelet decomposition (Mallat 1989; Daubechies 1992; Wang et al. 2018) as approximation $X_{A}$ and detail coefficients $X_{D_{i}},i\in\{1,2,\dots,f\}$ series corresponding to $m$ frequency levels from low to high.
@@
To demonstrate our proposed BEAT approach, specifically tailored for frequency-domain methods, we utilize WPMixer (Murad, Aktukmak, and Yilmaz 2025) as an instantiation. The approximation coefficient series at level $f$ is preserved, ensuring a concise and efficient representation.
@@
Frequency-Specific Monitor
The Frequency-Specific Monitor (FSM) is designed to track the training status of the network for each frequency, supporting the dynamical adjustment of the learning pace for these networks inspired by (Peng et al. 2022). To monitor the learning status for each frequency, it is necessary to compare their prediction with the corresponding targets. To this end, we take advantage of the ground-truth sequence $Y$ and utilize the wavelet decomposition to transform $Y$ in the temporal domain as the approximation coefficient series $Y_{A}$ and detail coefficient series $Y_{D_{i}}$ where $i\in\{1,2,\dots,f\}$ and $f$ is the wavelet decomposition level. The wavelet decomposition process can be formulated as:
@@
Dynamical Gradient Balancer
@@
Dataset # VAR Size FRQ
Weather 21 (36792, 5271, 10540) 10 min
ETTh1/h2 7 (8545, 2881, 2881) 1 hour
ETTm1/m2 7 (34465, 11521, 11521) 15 min
Traffic 862 (12185, 1757, 3509) 1 hour
ECL 321 (18317, 2633, 5261) 1 hour
@@
Table 1: Descriptions of datasets used for evaluation.
Experiments
Experimental Settings
Datasets.
We benchmark our method over seven challenging datasets in long-term series forecasting. (1) The weather dataset is collected every 10 minutes throughout 2020, containing 21 meteorological indicators such as air temperature and humidity. (2) ETT dataset (Zhou et al. 2021) contains data collected from electricity transformers, including load and oil temperature measurements recorded every 15 minutes from July 2016 to July 2018. There are four sub-datasets in ETT, including ETTh1, ETTh2, ETTm1, and ETTm2. (3) The traffic dataset comprises hourly data from the California Department of Transportation, detailing road occupancy rates measured by various sensors on San Francisco Bay Area freeways. (4) ECL dataset (Li et al. 2019) records hourly electricity consumption data from 321 clients. Table 1 provides data statistics of these datasets. Size denotes the number of time points in the (Train, Validation, Test) sets, respectively.
Baselines.
There are nine state-of-the-art baseline methods for long-term time series forecasting are compared, including WPMixer (Murad, Aktukmak, and Yilmaz 2025), TimeMixer (Wang et al. 2024a), iTransformer (Liu et al. 2024), TSMixer (Ekambaram et al. 2023), PatchTST (Nie et al. 2023), Crossformer (Zhang and Yan 2023), TiDE (Das et al. 2023), and DLinear (Zeng et al. 2023).
Index FSM DGB Weather ETTh1
MSE MAE MSE MAE
(1) 0.243 0.269 0.422 0.423
(2) 0.239 0.263 0.415 0.419
@@
Ours 2025 2024 2024 2023 2023 2023 2023 2023
96 0.160 0.198 0.162 0.204 0.163 0.209 0.174 0.214 0.175 0.247 0.186 0.227 0.195 0.271 0.202 0.261 0.195 0.252
192 0.207 0.241 0.209 0.246 0.208 0.250 0.221 0.254 0.224 0.294 0.234 0.265 0.209 0.277 0.242 0.298 0.237 0.295
336 0.258 0.279 0.263 0.287 0.251 0.287 0.278 0.296 0.262 0.326 0.284 0.301 0.273 0.332 0.287 0.335 0.282 0.331
720 0.329 0.332 0.339 0.338 0.339 0.341 0.358 0.347 0.349 0.348 0.356 0.349 0.379 0.401 0.351 0.386 0.345 0.382
Weather Avg 0.239 0.263 0.243 0.269 0.240 0.271 0.258 0.278 0.253 0.304 0.265 0.285 0.264 0.320 0.271 0.320 0.265 0.315
96 0.360 0.392 0.368 0.394 0.375 0.400 0.386 0.405 0.387 0.411 0.460 0.447 0.423 0.448 0.479 0.464 0.397 0.412
192 0.417 0.414 0.420 0.418 0.429 0.421 0.441 0.436 0.441 0.437 0.512 0.477 0.471 0.474 0.525 0.492 0.446 0.441
336 0.443 0.427 0.452 0.433 0.484 0.458 0.487 0.458 0.507 0.467 0.546 0.496 0.570 0.546 0.565 0.515 0.489 0.467
720 0.439 0.443 0.449 0.449 0.498 0.482 0.503 0.491 0.527 0.548 0.544 0.517 0.653 0.621 0.594 0.558 0.513 0.510
ETTh1 Avg 0.415 0.419 0.422 0.423 0.447 0.440 0.454 0.447 0.466 0.467 0.516 0.484 0.529 0.522 0.541 0.507 0.461 0.457
96 0.273 0.329 0.282 0.334 0.289 0.341 0.297 0.349 0.308 0.357 0.308 0.355 0.745 0.584 0.400 0.440 0.340 0.394
192 0.355 0.378 0.359 0.385 0.372 0.392 0.380 0.400 0.395 0.404 0.393 0.405 0.877 0.656 0.528 0.509 0.482 0.479
336 0.368 0.401 0.374 0.404 0.386 0.414 0.428 0.432 0.428 0.434 0.427 0.436 1.043 0.731 0.643 0.571 0.591 0.541
720 0.399 0.416 0.405 0.427 0.412 0.434 0.427 0.445 0.443 0.451 0.436 0.450 1.104 0.763 0.874 0.679 0.839 0.661
ETTh2 Avg 0.349 0.381 0.355 0.387 0.364 0.395 0.383 0.407 0.394 0.412 0.391 0.411 0.942 0.684 0.611 0.550 0.563 0.519
96 0.307 0.342 0.314 0.350 0.320 0.357 0.334 0.368 0.331 0.378 0.352 0.374 0.404 0.426 0.364 0.387 0.346 0.374
192 0.355 0.369 0.358 0.375 0.361 0.381 0.377 0.391 0.386 0.399 0.390 0.393 0.450 0.451 0.398 0.404 0.382 0.391
336 0.378 0.390 0.384 0.395 0.390 0.404 0.426 0.420 0.426 0.421 0.421 0.414 0.532 0.515 0.428 0.425 0.415 0.415
720 0.445 0.427 0.448 0.432 0.454 0.441 0.491 0.459 0.489 0.465 0.462 0.449 0.666 0.589 0.487 0.461 0.473 0.451
ETTm1 Avg 0.371 0.382 0.376 0.388 0.381 0.395 0.407 0.410 0.408 0.416 0.406 0.407 0.513 0.495 0.419 0.419 0.404 0.408
96 0.164 0.245 0.171 0.253 0.175 0.258 0.180 0.264 0.179 0.282 0.183 0.270 0.287 0.366 0.207 0.305 0.193 0.293
192 0.228 0.290 0.234 0.294 0.237 0.299 0.250 0.309 0.244 0.305 0.255 0.314 0.414 0.492 0.290 0.364 0.284 0.361
336 0.285 0.327 0.292 0.333 0.298 0.340 0.311 0.348 0.320 0.357 0.309 0.347 0.597 0.542 0.377 0.422 0.382 0.429
720 0.384 0.386 0.387 0.390 0.391 0.396 0.412 0.407 0.419 0.432 0.412 0.404 1.730 1.042 0.558 0.524 0.558 0.525
ETTm2 Avg 0.265 0.312 0.271 0.317 0.275 0.323 0.288 0.332 0.290 0.344 0.290 0.334 0.757 0.610 0.358 0.404 0.354 0.402
96 0.149 0.240 0.150 0.241 0.153 0.247 0.148 0.240 0.190 0.299 0.190 0.296 0.219 0.314 0.237 0.329 0.210 0.302
192 0.160 0.251 0.162 0.252 0.166 0.256 0.162 0.253 0.216 0.323 0.199 0.304 0.231 0.322 0.236 0.330 0.210 0.305
336 0.178 0.268 0.179 0.270 0.185 0.277 0.178 0.269 0.226 0.334 0.217 0.319 0.246 0.337 0.249 0.344 0.223 0.319
720 0.215 0.299 0.217 0.304 0.225 0.310 0.225 0.317 0.250 0.353 0.258 0.352 0.280 0.363 0.284 0.373 0.258 0.350
ECL Avg 0.176 0.265 0.177 0.267 0.182 0.272 0.178 0.270 0.220 0.327 0.216 0.318 0.244 0.334 0.251 0.344 0.225 0.319
96 0.459 0.278 0.465 0.286 0.462 0.285 0.395 0.268 0.499 0.344 0.526 0.347 0.644 0.429 0.805 0.493 0.650 0.396
192 0.470 0.282 0.475 0.290 0.473 0.296 0.417 0.276 0.540 0.370 0.522 0.332 0.665 0.431 0.756 0.474 0.598 0.370
336 0.487 0.294 0.489 0.296 0.498 0.296 0.433 0.283 0.557 0.378 0.517 0.334 0.674 0.420 0.762 0.477 0.605 0.373
720 0.510 0.312 0.527 0.318 0.506 0.313 0.467 0.302 0.586 0.397 0.552 0.352 0.683 0.424 0.719 0.449 0.645 0.394
Traffic Avg 0.482 0.292 0.489 0.297 0.484 0.297 0.428 0.282 0.546 0.372 0.529 0.341 0.667 0.426 0.760 0.473 0.625 0.383
1st Count: 28 30 0 0 1 0 7 6 0 0 0 0 0 0 0 0 0 0
[…]
Long-Term Time Series Forecasting Results
@@
Table 4: Ablation study of the wavelet type’s design choice.
Index
Levels of
Decomposition $f$
Weather ETTh1
@@
Table 5: Ablation study of levels of decomposition in the wavelet transform.
Ablation Study
@@
Index
Discrepancy
Calculation
Weather ETTh1
MSE MAE MSE MAE
(1) RMSE 0.244 0.268 0.421 0.425
(2) R-Squared 0.241 0.265 0.418 0.423
(3) MSE 0.239 0.263 0.415 0.419
(4) MAE 0.240 0.264 0.416 0.419
Table 6: Ablation study of the design for discrepancy calculation.
Index Modulation Weather ETTh1
MSE MAE MSE MAE
(1) Loss 0.242 0.268 0.420 0.421
(2) Gradient 0.239 0.263 0.415 0.419
Table 7: Ablation study of the design for learning process modulation.
@@
To evaluate the effect of the discrepancy calculation metric used in Equations 2 and 3, we conduct experiments on the metric for measuring the discrepancy between the predicted coefficients and the ground-truth ones, reflecting the learning process of networks for different frequencies. As shown in Table 6, using MSE for calculating the discrepancy can achieve the best performance against other metrics, including RMSE, R-Squared, and MAE. MSE outperforms other metrics because by squaring the errors, it disproportionately penalizes larger discrepancies, leading to a more accurate measurement.